# MROVSeg: Breaking the Resolution Curse of Vision-Language Models in Open-Vocabulary Semantic Segmentation

### v1

@@ 4 Experiments
Table 1: Performance comparison with state-of-the-art methods. †: We cite the reproduction performance of these methods[35, 13] trained with full COCO Stuff dataset from previous works[36, 11]. PC: Pascal Context, VOC: Pascal VOC.
Method VL-Model Training Dataset Extra Backbone ADE-847 PC-459 ADE-150 PC-59 VOC-20
@@
MROVSeg CLIP ViT-B/16 COCO Stuff - 12.1 19.6 32.0 58.5 95.5
@@
MROVSeg CLIP ViT-L/14 COCO Stuff - 16.1 24.1 36.9 64.1 97.6
4.1 Setting
We train our models on COCO-Stuff [3] dataset which comprises 164K images with densely annotated masks spanning 171 categories. We first evaluate MROVSeg on five well-established mainstream open-vocabulary semantic segmentation benchmarks for standard evaluation. We further evaluate MROVSeg on Cityscapes [12] benchmarks to explore the ability of handling high-resolution image input. We follow common practice [35, 2] to measure the segmentation performance by mean intersection over union (mIoU) score.
Datasets. The standard benchmarks contains three dataset: ADE [43], Pascal Context [26], and Pascal VOC [15]. The ADE dataset contains around 20K and 2K images for training and validation, respectively. This dataset is annotated with 150 and 847 categories, resulting in two separate segmentation benchmarks, namely ADE-150 and ADE-847. Similarly, the Pascal Context dataset has 5K images for both training and validation. It is annotated with 59 and 459 classes, forming two benchmarks known as PC-59 and PC-459. The Pascal VOC dataset comprises 1464 and 1449 images for training and validation, encompassing annotated masks across 20 semantic categories.
Implementation Details. We adopt the vanilla ViT block as transformer block in Multi-Res Adapter, and we use 6 blocks with 12 attention heads, 100 query tokens, feature dimension is 768 as default. We choose OpenAI pretrained ViT-based CLIP [28] models in all experiments for better reproducibility. We empirically choose [CLS] token from CLIP layer 9, CLIP ViT-B/16 model, and subsequent layers for Multi-grained Masked Attention. i.e., we use last 3 blocks for Multi-grained Masked Attention. For CLIP ViT-L/14 model, we [CLS] token from CLIP layer 18, and use subsequent layers for Multi-grained Masked Attention. Our models are trained with image input resolution $640\times 640$ , and slice and downsample to $320\times 320$ to fit the CLIP input resolution. More implementation details are in Appendix Sec. B.
@@ 4.1 Setting
Figure 7: Effect of decoupled attention decoding for multi-grained semantics. With single attention mask decoding, the model is unable to recognize the class painting (h) in the wall even with proper predicted proposals (g) since the spatial cues are overwhelmed by background noise (i). Our decoupled attention mask decoding effectively split the global and local semantics in LR (c) and HR (f) attention masks, performing enhanced recognition (e).
4.2 Standard Evaluation
Comparisons with State-of-the-Art Methods. In Tab. 1, we compare MROVSeg with current state-of-the-art open-vocabulary semantic segmentation methods on standard benchmarks. First of all, our method significantly outperforms other state-of-the-art methods all various open-vocabulary semantic segmentation benchmarks. Specifically, our method supasses state-of-the-art method SAN [30] with the same CLIP backbones on all benchmarks by remarkable margins ($+1.4\%$ mIoU for ADE-847, $+5.9\%$ mIoU for PC-459, $+3.1\%$ mIoU for ADE-150, $+3.1\%$ mIoU for PC-59, and $+0.9$ for VOC-20 with CLIP ViT-B backbone, $+2.4\%$ mIoU for ADE-847, $+7.0\%$ mIoU for PC-459, $+3.6\%$ mIoU for ADE-150, $+3.9\%$ mIoU for PC-59, and $+2.1$ for VOC-20 with CLIP ViT-B backbone). Compared to methods with additional image backbones, our model outperforms EBSeg (with SAM) by $+0.9,+1.0,+3.2,+1.2,+1.1$ mIoUfor ADE-847, PC-459, ADE-150, PC-59, and VOC-20 respectively. In addition, our models outperforms other convolution-based high-resolution trained method FC-CLIP and SED. Fig. 8 shows the qualitative comparison between MROVSeg and state-of-the-art methods (SAN [36] and EBSeg [30]). Evidently, MROVSeg can segment objects more precisely (the first row, class sofa), and provide more detailed mask prediction (the second and third row, more accurate object boundaries).
Capability of Handling High-Rsolution Images. Fig.6 show the image size distribution of the evaluated in Tab.1. It is noteworthy that the size distribution of the test datasets are highly overlapped with COCO, where the image size primarily distributed from $[250,800]$. To evaluate the ability of handling high-resolution image input, we train MROVSeg on COCO-Stuff with $1024\times 1024$ image resolution and compare the performance scalability with convolution-based high-resolution pretrained method FC-CLIP on Cityscapes benchmark. Notably, with high resolution inputs ($2048\times 1024$), MROVSeg reaches comparable performance while scaling up the backbone, which greatly enhanced than other ViT-based methods (EBSeg and SAN). The size adaptability of ViTs have been demonstrated to be worse than ConvNets [38]. Conversely, with our specifically designed multi-resolution training, MROVSeg is able to match the performance of convolution-based method on high-resolution inputs.
Multi.Ada. Mul.Ma.Att. Hier.Dec. PC-59 A-150 A-847 PC-459 VOC-20
52.3 25.4 9.0 10.8 93.6
55.6 28.7 10.5 14.9 94.7
57.1 30.5 11.3 18.0 95.1
58.5 32.0 12.1 19.6 95.5
@@ 4.2 Standard Evaluation
00 6768 116.5 27.6 53.7 10.1 15.7 94.7
$1.0$ 13493 293.1 27.8 54.5 9.9 15.0 94.4
$0.75$ 16261 225.7 30.5 54.3 10.2 15.9 95.1
$0.625$ 12942 184.5 32.1 57.9 12.0 19.3 95.7
$0.5$ 9779 142.8 32.0 58.5 12.1 19.6 95.5
$0.25$ 7560 92.5 28.5 56.2 10.5 17.0 94.7
Table 4: Ablation study on various designs in Multi-Res Adapter. Multi-Res Adapter contains Multi-Res Fusion (MRF) module and ViT blocks. We first show the impact of spatial fusion method (a) and multi-resolution fusion (b). Then we give the impact of number of blocks (c), number of fusion CLIP layer (ViT-B model) (d), and the channel width (e). Finally, we show the impact of query number (f). The default setting are marked underline.
(a) Spatial Fusion A-847 PC-459 A-150 PC-59 VOC-20
Concat. 11.8 18.0 31.5 57.0 94.6
Depth.Conv. 12.1 19.6 32.0 58.5 95.5
(b) Multi-Res Fusion A-847 PC-459 A-150 PC-59 VOC-20
Add 11.7 18.4 31.6 56.9 94.5
Concat. 11.4 18.3 31.0 56.1 94.8
Scale.Attn. 12.1 19.6 32.0 58.5 95.5
(c) Block Num A-847 PC-459 A-150 PC-59 VOC-20
3 8.2 11.5 26.7 53.5 93.8
6 12.1 19.6 32.0 58.5 95.5
9 11.9 19.6 31.5 59.2 95.4
(d) Fusion Layer A-847 PC-459 A-150 PC-59 VOC-20
{3,6,9} 11.2 17.4 31.6 56.5 95.6
{3,6,9,12} 11.7 18.8 31.4 57.0 94.4
{stem,3,6,9} 11.8 19.4 32.1 58.1 94.9
{stem,3,6,9,12} 12.1 19.6 32.0 58.5 95.5
(e) Channel Width A-847 PC-459 A-150 PC-59 VOC-20
384 9.6 16.1 28.4 56.8 94.2
768 12.1 19.6 32.0 58.5 95.5
1024 12.0 19.1 32.4 58.0 95.7
(f) Query Num A-847 PC-459 A-150 PC-59 VOC-20
100 12.1 19.6 32.0 58.5 95.5
200 11.9 18.3 31.8 57.0 93.8
300 12.2 19.3 32.4 57.5 95.2
Figure 8: Qualitative comparison with SAN [36] and EBSeg [30].
4.3 Ablation Study
For all ablation experiments, we employ our CLIP ViT-B/16 based model as our ablated baseline.
Components Analysis. Tab. 2 shows the performance effect of the main components in MROVSeg. The baseline adopts single resolution CLIP plain ViT feature for mask decoding, and use the masked-attention for mask class recognition as in MaskCLIP and SAN. The baseline obtains the mIoU scores of 52.3%, 22.4%, 8.5%, 10.8% and 93.6% for PC-59, ADE-150, ADE-847, PC-459 and VOC-20. When we use multi-resolution images with High-Res Adapter, it has the mIoU scores of 55.6%, 28.7%, 10.5%, 16.8% and 94.7% for PC-59, ADE-150, ADE-847, PC-459 and VOC-20, which outperforms baseline by 3.3%, 6.3%, 2.0%, 6.0% and 1.1%. After use hierarchical features to decode mask, it obtains the mIoU scores of 57.1%, 30.5%, 11.3%, 14.9% and 95.1% for PC-59, ADE-150, ADE-847, PC-459 and VOC-20, which outperforms baseline by 4.8%, 5.1%, 2.3%, 7.2% and 1.5%. Finally we integrate the Multi-grained Masked-Attention to the model, it reaches 58.5%, 32.0%, 12.1%, 19.6% and 95.5% for PC-59, ADE-150, ADE-847, PC-459 and VOC-20, which outperforms baseline by 6.2%, 6.6%, 3.1%, 8.8% and 1.9%.
Single Resolution vs. Multi Resolution. Denote the slice ratio as $p=\frac{slice\;\,resolution}{input\;\,resolution}$, we examine the effect of $p$ value in Tab. 3. Notably, only using single resolution feature, i.e., $p=0$, or directly adopting high-resolution image as CLIP input, i.e., $p=1.0$ lead to significant performance degradation. While some $p$ values with overlapped slicing (such as $0.625$) obtains better performance on some datasets(such as ADE-150 and VOC-20), we choose the default $p$ value as $0.5$ considering computation overhead.
Micro-design Choice of Multi-Res Adapter. We conduct experiment to examine the different micro-design choice in Multi-Res Adapter in Tab. 3. We first perform experiments on MRF module. In (a), we examine different spatial restore strategies. The depthwise convolution is significantly better than concatenation. And we present the impact of local-global modeling methods (b), and the best results on all benchmarks are achieved by scale attention. Then for the ViT block in Multi-Res Adapter, we present the effect of block number in (c), we find that although increase the block number to 9 can bring slight improvement on some benchmarks (PC-59 and PC-459), it would incur more computation cost. In (d), we examine the effect of CLIP layers whose feature we adopt. In (e), we present the impact of channel width we use. Finally, we examine the effect of query number in (f).
@@ 4.3 Ablation Study
ZSSeg 530.8 22302.1 0.3 15h59min
OVSeg 532.6 19345.6 0.4 17h54min
CAT-Seg 433.7 2121.1 2.0 7h41min
FC-CLIP 221.3 680.0 2.3 2d5h
EBSeg 210.9 867.5 4.7 18h33min
@@
Computation Overhead. We compare the computation overhead of MROVSeg with recent methods [14, 22, 11, 30, 38] in Tab. 5. We measure the number of parameter, giga floating-point operations per second, inference FPS and training time. Our method exhibits strong efficiency among these methods in terms of both training and inference. This is achieved by 1) single CLIP backbone do not need additional image backbones; 2) slice then input strategy avoid quadratic computation cost with regard to input image size.

### v2

@@ 4 Experiments
Table 1: Semantic segmentation performance comparison with state-of-the-art methods. †: We cite the reproduction performance of these methods[35, 9] trained with full COCO Stuff dataset from previous works[36, 6]. PC: Pascal Context, VOC: Pascal VOC.
Method VL-Model Training Dataset End to End Training Extra Backbone ADE-847 PC-459 ADE-150 PC-59 VOC-20
@@
CAT-Seg [6] CLIP ViT-B/16 COCO Stuff - 12.0 19.0 31.8 57.5 94.6
MROVSeg CLIP ViT-B/16 COCO Stuff - 12.9 19.2 32.4 58.7 95.8
@@
CAT-Seg [6] CLIP ViT-L/14 COCO Stuff - 16.0 23.8 37.9 63.3 97.0
MAFT+ [15] CLIP ConvNext-L COCO Stuff - 15.1 21.6 36.1 59.4 96.5
MROVSeg CLIP ViT-L/14 COCO Stuff - 16.4 24.0 36.9 64.3 97.6
4.1 Settings
For open-vocabulary semantic segmentation task, We train our models on COCO-Stuff [2] dataset which comprises 164K images with densely annotated masks spanning 171 categories. Then we evaluate MROVSeg on five well-established open-vocabulary semantic segmentation benchmarks for standard evaluation. We further evaluate MROVSeg on Cityscapes [7] benchmarks to explore the ability of handling high-resolution image input. We follow common practice [35, 1] to measure the segmentation performance by mean intersection over union (mIoU) score. For open-vocabulary panoptic segmentation task, we train MROVSeg on COCO-Panoptic [20] dataset. Then we evaluate zero-shot performance of MROVSeg on ADE [41] panoptic benchmark, and measure the panoptic segmentation performance by panoptic quality (PQ), segmentation qualitiy (SQ) and recognition quality (RQ).
Datasets The standard semantic segmentation benchmarks contains three dataset: ADE [41], Pascal Context [23], and Pascal VOC [11]. The ADE dataset contains around 20K and 2K images for training and validation, respectively. This dataset is annotated with 150 and 847 categories, resulting in two separate segmentation benchmarks, namely ADE-150 and ADE-847. Similarly, the Pascal Context dataset has 5K images for both training and validation. It is annotated with 59 and 459 classes, forming two benchmarks known as PC-59 and PC-459. The Pascal VOC dataset comprises 1464 and 1449 images for training and validation, encompassing annotated masks across 20 semantic categories.
Table 2: Panoptic segmentation performance comparison with state-of-the-art methods.
ADE COCO
open-vocabulary close-set
Method PQ SQ RQ PQ SQ RQ
FreeSeg [25] 16.3 - - - -
MaskCLIP [10] 15.1 70.4 19.2 - - -
ODISE [34] 22.6 - - - -
OPSNet [5] 19.0 52.4 23.0 - - -
FCCLIP [39] 26.8 71.5 32.2 54.4 44.6 63.7
MAFT-Plus [15] 27.1 73.5 32.9 - - -
MROVSeg 27.3 72.8 33.4 52.0 41.4 60.9
Implementation Details We adopt the vanilla ViT block as transformer block in Multi-Res Adapter, and we use 6 blocks with 12 attention heads, 100 query tokens, feature dimension is 768 as default. We choose OpenAI pretrained ViT-based CLIP [26] models in all experiments for better reproducibility. We empirically choose [CLS] token from CLIP layer 9, CLIP ViT-B/16 model, and subsequent layers for Multi-grained Masked Attention. i.e., we use last 3 blocks for Multi-grained Masked Attention. For CLIP ViT-L/14 model, we [CLS] token from CLIP layer 18, and use subsequent layers for Multi-grained Masked Attention. Our models are trained with image input resolution $640\times 640$ , and slice and downsample to $320\times 320$ to fit the CLIP input resolution. More implementation details are in Supplementary Material.
@@ 4.1 Settings
4.2 Accuracy Evaluation
Open-Vocabulary Semantic Segmentation We compare the semantic segmentation performance of MROVSeg with current state-of-the-art methods in Tab. 1. First of all, our method significantly outperforms other state-of-the-art methods all various on most open-vocabulary semantic segmentation benchmarks. Specifically, our method supasses state-of-the-art method CAT-Seg [28] with the same CLIP backbones on four benchmarks by remarkable margins ($+0.9\%$ mIoU for ADE-847, $+0.2\%$ mIoU for PC-459, $+0.6\%$ mIoU for ADE-150, $+1.2\%$ mIoU for PC-59, and $+1.2$ for VOC-20 with CLIP ViT-B backbone, $+0.4\%$ mIoU for ADE-847, $+0.2\%$ mIoU for PC-459, $+1.0\%$ mIoU for PC-59, and $+0.6$ for VOC-20 with CLIP ViT-B backbone). Compared to methods with additional image backbones, our model outperforms EBSeg [28](with SAM) by $+2.7,+3.0,+4.1,+4.1,+1.4$ mIoUfor ADE-847, PC-459, ADE-150, PC-59, and VOC-20 respectively. In addition, our models outperforms other convolution-based high-resolution trained method FC-CLIP [39] and SED [33]. Fig. 8 shows the qualitative comparison between MROVSeg and state-of-the-art methods (SAN [36] and EBSeg [28]). Evidently, MROVSeg can segment objects more precisely (the first row, class sofa), and provide more detailed mask prediction (the second and third row, more accurate object boundaries).
The size adaptability of ViTs have been demonstrated [39] to be worse than ConvNets. Fig. 7 shows that the image sizes of the datasets evaluated in Tab. 1 are primarily distributed from $[250,800]$. To evaluate the ability of handling high-resolution image input, we retrain MROVSeg on COCO-Stuff with $1024\times 1024$ resolution evaluate performance on Cityscapes benchmark. Notably, MROVSeg reaches comparable performance to convolution-based methods [39] while scaling up the backbone, which greatly enhanced than other ViT-based methods (EBSeg [28] and SAN [36]).
Figure 8: Qualitative comparison with SAN [36] and EBSeg [28].
@@ 4.2 Accuracy Evaluation
00 6768 116.5 28.6 54.7 10.6 16.4 94.5
$1.0$ 13493 293.1 28.8 54.5 10.5 15.6 94.4
$0.75$ 16261 225.7 31.5 54.0 11.5 15.9 95.1
$0.625$ 12942 184.5 32.1 59.1 12.0 19.6 95.6
$0.5$ 9779 142.8 32.4 58.7 12.9 19.2 95.8
$0.25$ 7560 92.5 27.5 55.7 9.9 16.0 93.7
@@
ZSSeg [35] 530.8 22302.1 0.3 15h59min
OVSeg [19] 532.6 19345.6 0.4 17h54min
CAT-Seg [6] 433.7 2121.1 2.0 7h41min
FC-CLIP [39] 221.3 680.0 2.3 2d5h
EBSeg [28] 210.9 867.5 4.7 18h33min
@@
Open-Vocabulary Panoptic Segmentation In Tab. 2, We evalute the panoptic segmentation performance of MROVSeg on mainstream open-vocabulary panoptic segmentation benchmarks ADE [41]. It is noteworthy that our method surpass previous arts on panoptic quality (PQ) and recognition quality (RQ), achieving new state-of-art open-vocabulary panoptic segmentation performance. Furthermore, close-set panoptic segmentation results indicate that training MROVSeg model does not lead to severe overfitting issue on the base dataset.
4.3 Efficiency Evaluation
Memory Consumption Denote the slice ratio as $p=\frac{slice\;\,resolution}{input\;\,resolution}$, we examine the effect of $p$ value in Tab. 3. Notably, only using single resolution feature, i.e., $p=0$, or directly adopting high-resolution image as CLIP input, i.e., $p=1.0$ lead to significant performance degradation. While some $p$ values with overlapped slicing (such as $0.625$) obtains better performance on some datasets(such as ADE-150 and VOC-20), we choose the default $p$ value as $0.5$ considering computation overhead.
@@ 4.3 Efficiency Evaluation
Computation Overhead We compare the computation overhead of MROVSeg with recent methods [10, 19, 6, 28, 39] in Tab. 4. We measure the number of parameter, giga floating-point operations per second, inference FPS and training time. Our method exhibits strong efficiency among these methods in terms of both training and inference. This is achieved by 1) single CLIP backbone do not need additional image backbones; 2) slice then input strategy avoid quadratic computation cost with regard to input image size. More detailed parameter settings are in Supplementary Material.
Method PC-59 A-150 A-847 PC-459 VOC-20
Baseline 52.3 25.4 9.0 10.8 92.9
+ High-res Features 54.1 28.1 10.4 13.0 93.6
+ Multi-Res Adapter 55.6 28.7 10.5 14.9 94.7
+ Masked Attention 56.4 30.9 11.8 17.9 95.5
+ Multi-grained Masked Attention 58.7 32.4 12.9 19.2 95.8
Table 6: Effect of Hierarchical Mask Decoding and Image-conditioned Text Feature.
Method PC-59 A-150 A-847 PC-459 VOC-20
MROVSeg w/o Hier. Mask Dec. 57.1 30.5 11.3 18.0 95.1
MROVSeg w/o Img-cond. Text Feat. 58.5 32.0 12.1 19.6 95.5
MROVSeg 58.7 32.4 12.9 19.2 95.8
4.4 Ablation Study
Table 7: Ablation study on various designs in Multi-Res Adapter. The default setting are marked underline.
(a) Spatial Fusion A-847 PC-459 A-150 PC-59 VOC-20
Concat. 12.0 18.4 31.3 57.1 94.8
Depth.Conv. 12.9 19.2 32.4 58.7 95.8
(b) Multi-Res Fusion A-847 PC-459 A-150 PC-59 VOC-20
Add 12.0 17.7 31.2 55.9 94.5
Concat. 12.4 19.8 31.8 58.7 95.6
Scale.Attn. 12.9 19.2 32.4 58.7 95.8
(c) Block Num A-847 PC-459 A-150 PC-59 VOC-20
3 9.8 12.5 27.7 53.9 94.1
6 12.9 19.2 32.4 58.7 95.8
9 12.7 19.3 31.9 59.2 95.4
(d) Fusion Layer A-847 PC-459 A-150 PC-59 VOC-20
{3,6,9} 11.8 19.1 31.6 56.5 95.9
{3,6,9,12} 12.4 19.8 32.0 58.0 95.4
{stem,3,6,9} 12.8 19.4 32.6 58.1 94.9
{stem,3,6,9,12} 12.9 19.2 32.4 58.7 95.8
(e) Channel Width A-847 PC-459 A-150 PC-59 VOC-20
384 10.4 17.0 27.5 56.1 94.2
768 12.9 19.2 32.4 58.7 95.8
1024 11.1 19.7 31.0 59.5 96.1
(f) Query Num A-847 PC-459 A-150 PC-59 VOC-20
100 12.9 19.2 32.4 58.7 95.8
200 11.9 18.3 31.8 57.0 93.8
300 12.2 18.7 32.4 57.5 95.2
For ablation experiments, except for Tab. 8, we all employ our CLIP ViT-B/16 based model as our ablated baseline.
Components Analysis Tab. 5 shows the performance effect of the main components in MROVSeg. The baseline adopts single resolution CLIP plain ViT feature for mask decoding, and uses the MaskPooling [34] for mask class recognition. The baseline obtains the mIoU scores of 52.3%, 22.4%, 8.5%, 10.8% and 93.6% for PC-59, ADE-150, ADE-847, PC-459 and VOC-20. We first introduce multi-resolution CLIP features, which marginally outperforms baseline. Then we introduce the Multi-Res Adapter, the model significantly outperforms baseline by 4.8%, 5.1%, 2.3%, 7.2% and 1.5%. Next we integrate the Masked-Attention [36] to the model, it slightly outperforms MaskPooling. Finally we integrate Multi-grained Masked Attention, the model performance reaches 58.7%, 32.4%, 12.9%, 19.2% and 95.8% for PC-59, ADE-150, ADE-847, PC-459 and VOC-20, which outperforms baseline by 6.4%, 7.0%, 3.9%, 8.4% and 2.2%. Futhermore, we show the effect of of Hierarchical Mask Decoding and Image-conditioned Text Feature in Tab. 6, both showing consistent improvements.
Effect of Multi-Res Adapter We first conduct experiment to examine the different micro-design choice in Multi-Res Adapter in Tab. 3. In (a), we examine different spatial restore strategies in MRF module. The depthwise convolution is significantly better than concatenation. And we present the impact of local-global modeling methods (b), and the best results on all benchmarks are achieved by scale attention. Then for the ViT blocks setting, we present the effect of block number in (c). Increasing the block number to 9 brings limited improvement while incurring heavy computation cost. In (d), we examine the effect of CLIP layers whose feature we adopt. In (e), we present the impact of channel width we use. Finally, we examine the effect of query number in (f). We further compare the performance of MROVSeg with ViT-based and convolution-based CLIP models in Tab. 8. The results indicate that ViT-based encoder with a sliding window approach outperforms convolution-based counterpart.
Figure 9: Effects of decoupled attention decoding for multi-grained semantics. The t-SNE [30] visualization shows that decoupled attention masks can further split the query embedding from different classes, making clear classification boundaries.
Table 8: Comparison between different VLMs. All Multi-grained Masked Attention is disabled to adapt ConvNext.
Method VLM PC-59 A-150 A-847 PC-459 VOC-20
MROVSeg ConvNext-L 57.0 28.4 10.1 15.8 93.3
MROVSeg ViT-L 58.5 30.4 13.4 20.1 95.4
Effect of Multi-grained Masked Attention As mentioned before, Tab. 5 and Fig. 4 quantitatively and qualitatively show the effectiveness of Multi-grained Masked Attention respectively. We further visualize the query [CLS] embeddings ( i.e., $\mathbf{X}_{\texttt{prop}}$) by t-SNE [30] dimensionality reduction within ADE-150 [41] benchmarks in Fig. 9. We color the embeddings based on the Hungarian matching with groundtruth. In (a), we can observe that the queries from the same classes are well-posed as masked attention can aggregate semantics. Conversely, with decoupled attention masks, the query embeddings from different classes are split further in (b), indicating the effectiveness of multi-grained masked attention in mask classification.